OpenStudy (anonymous):

can some one help me with the steps to get the answers of this problem? Find the unknown number. Three added to 4 times a number is equal to 5 less than 6 times the number. Find the unknown number. Three added to 4 times a number is equal to 5 less than 6 times the number. Find the unknown number. Three added to 4 times a number is equal to 5 less than 6 times the number.

OpenStudy (study_together):

Maybe this will help for #1: 3+4x=6x-5 Find x for the answer

OpenStudy (anonymous):

Three added to 4 times a number "x" is equal to 5 less than 6 times the number. 4x + 3 = 6x -5 8 = 2x 4 = x